How many cups of grape juice must be added to 40 cups of punch that is 5% grape juice to obtain a punch that is 10% grape juice?
Answer by scott8148(6628)

40(5%) + x(100%) = (40+x)(10%)

dividing by 5% ___ 40 + 20x = 80+2x

18x = 40